# DARNet: Dual Attention Refinement Network with Spatiotemporal Construction for Auditory Attention Detection

### v1

@@ Abstract
At a cocktail party, humans exhibit an impressive ability to direct their attention. The auditory attention detection (AAD) approach seeks to identify the attended speaker by analyzing brain signals, such as EEG signals. However, current AAD algorithms overlook the spatial distribution information within EEG signals and lack the ability to capture long-range latent dependencies, limiting the model’s ability to decode brain activity. To address these issues, this paper proposes a dual attention refinement network with spatiotemporal construction for AAD, named DARNet, which consists of the spatiotemporal construction module, dual attention refinement module, and feature fusion & classifier module. Specifically, the spatiotemporal construction module aims to construct more expressive spatiotemporal feature representations, by capturing the spatial distribution characteristics of EEG signals. The dual attention refinement module aims to extract different levels of temporal patterns in EEG signals and enhance the model’s ability to capture long-range latent dependencies. The feature fusion & classifier module aims to aggregate temporal patterns and dependencies from different levels and obtain the final classification results. The experimental results indicate that compared to the state-of-the-art models, DARNet achieves an average classification accuracy improvement of 5.9% for 0.1s, 4.6% for 1s, and 3.9% for 2s on the DTU dataset. While maintaining excellent classification performance, DARNet significantly reduces the number of required parameters. Compared to the state-of-the-art models, DARNet reduces the parameter count by 91%. Code is available at: https://github.com/fchest/DARNet.git.
@@ 1 Introduction
  • The DARNet achieves remarkable decoding accuracy within very short decision windows, surpassing the current state-of-the-art (SOTA) model by 5.9% on the DTU dataset and 4.9% on the KUL dataset, all under a 0.1-second decision window. Furthermore, compared to the current state-of-the-art model with 0.91 million training parameters, DARNet achieves further parameter reduction, requiring only 0.78 million parameters.
@@ 4.2 Ablation Study
Figure 2: AAD accuracy(%) of DARNet across all subjects on KUL, DTU, and MM-AAD datasets.
@@ 5.2 Ablation Analysis
5.3 Computational Cost
We compare the training parameter counts of our DARNet, SSF-CNN [37], MBSSFCC [15], and DBPNet [20], with the results shown in Table 4. The parameter count of DARNet is 51.6 times lower than that of SSF-CNN, 1331.5 times lower than that of MBSSFCC, and 0.91 times lower than that of DBPNet. Compared to other models, DARNet demonstrates superior parameter efficiency. Despite having fewer parameters, DARNet maintains good performance, indicating its ability to be applied in resource-constrained environments for AAD analysis, thus demonstrating practical utility.
Table 4: The training parameter counts of our DARNet and three open-source models. "M" denotes a million.

### v2

@@ Abstract
At a cocktail party, humans exhibit an impressive ability to direct their attention. The auditory attention detection (AAD) approach seeks to identify the attended speaker by analyzing brain signals, such as EEG signals. However, current AAD algorithms overlook the spatial distribution information within EEG signals and lack the ability to capture long-range latent dependencies, limiting the model’s ability to decode brain activity. To address these issues, this paper proposes a dual attention refinement network with spatiotemporal construction for AAD, named DARNet, which consists of the spatiotemporal construction module, dual attention refinement module, and feature fusion & classifier module. Specifically, the spatiotemporal construction module aims to construct more expressive spatiotemporal feature representations, by capturing the spatial distribution characteristics of EEG signals. The dual attention refinement module aims to extract different levels of temporal patterns in EEG signals and enhance the model’s ability to capture long-range latent dependencies. The feature fusion & classifier module aims to aggregate temporal patterns and dependencies from different levels and obtain the final classification results. The experimental results indicate that DARNet achieved excellent classification performance, particularly under short decision windows. While maintaining excellent classification performance, DARNet significantly reduces the number of required parameters. Compared to the state-of-the-art models, DARNet reduces the parameter count by 91%. Code is available at: https://github.com/fchest/DARNet.git.
@@ 1 Introduction
  • The DARNet achieves remarkable decoding accuracy within very short decision windows, surpassing the current state-of-the-art (SOTA) model by 5.9% on the DTU dataset and 4.9% on the KUL dataset, all under a 0.1-second decision window. Furthermore, compared to the current state-of-the-art model with 0.91 million training parameters, DARNet achieves further parameter reduction, requiring only 0.08 million parameters.
@@ 4.2 Ablation Study
Figure 2: AAD accuracy(%) of DARNet across all subjects on three datasets.
@@
4.3 Experimental Correction
To ensure fairness in comparisons, we aligned our previous experimental setup and data processing methods with those of DBPNet. However, during the code review for the final version of this paper, we found that DBPNet applies CSP to the data prior to dataset splitting, which can lead to data leakage. Consequently, we corrected the processing steps and conducted additional experiments, as shown in Table 4. The data preprocessing steps for the supplementary experiments are as follows:
  1. 1.
    Split each trial of the dataset into the first 90% for training and the last 10% for testing.
  2. 2.
    Fit the CSP transformation matrix using the training data, project the data to extract features, and apply the matrix to the testing data.
  3. 3.
    Apply a sliding window with 50% overlap to the processed training data, then randomly split into a training (90%) and a validation (10%) set. Similarly, apply a sliding window to the processed testing data as the final test set.
Following these steps, we rigorously avoided risk of data leakage. Experimental results show that both corrected DBPNet2 and DARNet2 exhibit a performance decline compared to the original DBPNet1 and DARNet1. However, DARNet2 still achieves state-of-the-art performance under most test conditions, as shown in Table 4. For instance, on the KUL dataset with a 0.1-second decision window, DARNet2 maintains a decoding accuracy of 89.2%.
Table 4: Experimental Correction on KUL, DTU, and MM-AAD dataset. The experimental setup for the results marked with 1 is consistent with DBPNet[20], where CSP is applied prior to dataset splitting. In contrast, for the results marked with 2, CSP is applied after dataset splitting.
Dataset Scene Model Decision Window
0.1-second 1-second 2-second
KUL audio-only DBPNet1 87.1 $\pm$ 6.55 95.0 $\pm$ 4.16 96.5 $\pm$ 3.50
DBPNet2 85.3 $\pm$ 6.22 94.4 $\pm$ 4.62 95.3 $\pm$ 4.63
DARNet1 91.6 $\pm$ 4.83 96.2 $\pm$ 3.04 97.2 $\pm$ 2.50
DARNet2 89.2 $\pm$ 5.50 94.8 $\pm$ 4.53 95.5 $\pm$ 4.89
DTU audio-only DBPNet1 75.1 $\pm$ 4.87 83.9 $\pm$ 5.95 86.5 $\pm$ 5.34
DBPNet2 74.0 $\pm$ 5.20 79.8 $\pm$ 6.91 80.2 $\pm$ 6.79
DARNet1 79.5 $\pm$ 5.84 87.8 $\pm$ 6.02 89.9 $\pm$ 5.03
DARNet2 74.6 $\pm$ 6.09 80.1 $\pm$ 6.85 81.2 $\pm$ 6.34
MM-AAD audio-only DBPNet1 91.4 $\pm$ 4.63 92.0 $\pm$ 5.42 92.5 $\pm$ 4.59
DBPNet2 90.0 $\pm$ 5.51 90.7 $\pm$ 5.68 91.6 $\pm$ 4.82
DARNet1 94.9 $\pm$ 4.79 96.0 $\pm$ 4.00 96.5 $\pm$ 3.59
DARNet2 91.5 $\pm$ 5.27 92.2 $\pm$ 4.54 92.8 $\pm$ 5.22
audio-visual DBPNet1 92.1 $\pm$ 4.47 92.8 $\pm$ 5.94 93.4 $\pm$ 4.86
DBPNet2 92.0 $\pm$ 5.51 93.0 $\pm$ 5.19 92.7 $\pm$ 6.04
DARNet1 95.8 $\pm$ 4.04 96.4$\pm$ 3.72 96.8 $\pm$ 3.44
DARNet2 92.7 $\pm$ 5.34 93.4$\pm$ 5.23 94.2 $\pm$ 4.84
@@ 5.2 Ablation Analysis
Table 5: The training parameter counts comparison. "M" denotes a million.
@@
5.3 Computational Cost
We compare the training parameter counts of our DARNet, SSF-CNN [37], MBSSFCC [15], and DBPNet [20], with the results shown in Table 5. The parameter count of DARNet is 51.6 times lower than that of SSF-CNN, 1331.5 times lower than that of MBSSFCC, and 10.4 times lower than that of DBPNet. Compared to other models, DARNet demonstrates superior parameter efficiency. Despite having fewer parameters, DARNet maintains good performance, indicating its ability to be applied in resource-constrained environments for AAD analysis, thus demonstrating practical utility.